1/2

Ирина Саверьевна Алексеева ,  Ирина Саверьевна Алексеева ,  Ирина Саверьевна Алексеева ,
Ode a Samarcanda  Odo al Samarkando  Ode to Samarkand
translated by Carlo Minnaja  translated by Vladimir Masalkin  translated by Malcolm Jones

Argiope information: The view of this version is restricted. Presumably there are copyright issues.

Mi kantas vin, ho, Samarkand’ Kaj laŭdas mi aspekton sunan! Kaj vian koron — Registan’ — Mi ĉiam pretas nomi juna.

I sing of you, oh, Samarkand’!

Your wondrous features win my praises. And your dear heart - the Registan - seems young to me and never ages.

Korbaton ĝian sub stelar’ Eksentis Ulughbek junece. Kaj daŭras de Timur centjar’ Kaj pli ĝi longas majestece!

Its steady beating Ulughbek

could sense beneath the constellations. Oh, Samarkand! Immortal pick of mighty Tamerlan’s creations.

Madrasoj sanktaj estas nun La ejoj por la art’ kaj sentoj. Kaj de belec’ ilia sun’ Centjarojn igis la momentoj.

Madrasses are your sacred schools, which cultivate the arts and learning. Their beauty overcomes time’s rules by centuries into moments turning.

La poezi’ de Navoi, Je amo kaj je forto riĉa, Tre helpis al popol’ de vi Fiera resti kaj feliĉa.

Here the great Navoi’s verse, so full of love and strong emotion, is cherished by your populace, made happy by its inspiration.

De minaretoj kupolar’ Similas al stelar’ ĉiela, Beleco de la ornamar’, Lazuro pura kaj tre hela.

The domes of minarets stand proud - like day-time stars in clear sky shining: so pure the blue, without a cloud and fine the spires in art combining.

…  …  …

2/2

En saĝa urbo de l’anim’ Moske’ najbaras kun preĝejo. Por homoj, kiuj amas vin, Vi estas bela rifuĝejo!

Oh, Samarkand! You’re wise through age,

with mosques and churches closely neighboured.

For those who love you, the world’s stage has no town better nor more favoured.

Ho, Samarkand’, disfloru pli Kaj gloru vin bonagoj fame! Kaj flugu al eterno vi, Restante apud ni ĉiame.

Oh, Samarkand! Live long and thrive, be glorified by each good action! And while eternity’s your prize, be here our lasting benefaction.

Translation of the Russian poem “Oда Самарканду” by Ирина Саверьевна Алексеева into itala by Carlo Minnaja (∗1940-03-19).

Arg-860-2206 (2014-11-13 12:31:20)

Tiu ĉi poem-italigo aperis en 2013 en la libro “Oда Самарканду”. Ода Самарканду: пер. на 62 языка / И. С. Алексеева ; ред.-сост.: А. Ионесов . – Москва. : Изд-во журн. “Юность”, 2013 . – 91 с. : ил., 2 л. ил. Pri la tradukinto vidu la retejon http: // it. wikipedia. org/ wiki/ Carlo_Minnaja .

Translation of the Russian poem “Oда Самарканду” by Ирина Саверьевна Алексеева into Esperanto by Vladimir Masalkin.

Arg-860-1734 (2013-03-24 15:52:41)

Translation of the Russian poem “Oда Самарканду” by Ирина Саверьевна Алексеева into English by Malcolm Jones.

Arg-860-2037 (2014-10-16 20:52:12)

Tiu ĉi poem-angligo aperis en 2013 en libro “Oда Самарканду” (= Odo por Samarkando). Sinjoro Malcolm Jones, kiu angligis la poemon, estas brita esperantisto. Lia adreso troviĝas en la UEAjarlibro de la jaro 2014 sur paĝo 126.

http://www.poezio.net/version?poem-id=860&version-id=2206,1734,2037